# ManiFeel: Benchmarking and Understanding Visuotactile Manipulation Policy Learning

### v1

@@ Appendix B Additional Details on Task Environments
Tactile Sorting (Real). The robot must classify and sort two visually similar spherical objects, a golf ball and a table tennis ball, into red and yellow bowls, respectively. The task is conducted under either normal or dim lighting conditions. Visual observations are captured using a front-facing Intel RealSense camera [^5] https://www.intelrealsense.com/depth-camera-d435/, while tactile feedback is provided by a GelSight Mini sensor mounted on the robot’s gripper. The two balls are placed on a ball stand. At the start of every new demonstration or evaluation rollout, the positions of the two balls are swapped to introduce the variation. OSC-yaw controller is used to control end-effector position and yaw rotation, while roll and pitch are kept fixed throughout the execution.
Tactile Sorting (Simulation). This task is conducted under dim lighting conditions to simulate degraded visual perception. A golf ball and a table tennis ball are placed on a ball stand whose position is randomly initialized within a range of $(-5\,\mathrm{cm},+5\,\mathrm{cm})$ along both $x$ and $y$ axes from its nominal position at the start of each demonstration or evaluation rollout. To increase task diversity, the positions of the balls (i.e., golf vs. table tennis) are also swapped between different episodes.
Tactile Exploration (Simulation). This task is performed in a vision-occluded environment, with the object (a cube-shaped battery) placed inside a closed container. At the beginning of each demonstration or evaluation, the object’s position is randomly placed at two predefined locations along the $x$-axis. Due to the limited and occluded visual input, the robot must rely solely on tactile feedback for object exploration, localization, and grasping.
Blind Insertion (Simulation). In this vision-occluded task, the socket’s position remains fixed throughout all rollouts, while the peg’s initial in-hand orientation is randomized up to $45^{\circ}$. This variation necessitates that the policy infer and adjust the in-hand pose through tactile sensing to ensure successful insertion. In this task, the gripper remains closed to ensure a stable hold on the peg during insertion.
@@
Insertion Tasks (Simulation). For USB and power plug insertion tasks, both the socket location and the in-hand orientation of the object (USB or plug) are randomized at the start of each rollout. The socket position is sampled within a $(-5\,\mathrm{cm},+5\,\mathrm{cm})$ range along the $x$ and $y$ axes, and the object’s orientation is randomized within $45^{\circ}$ of its nominal pose. In this task, the gripper remains closed to ensure a stable hold on the peg during insertion.
Gear Assembly (Simulation). In this industrial-style assembly task, the gear base location on the table and the in-hand orientation of the gear are both randomized at the start of each rollout. The gear base is perturbed within a $(-5\,\mathrm{cm},+5\,\mathrm{cm})$ range in the $x$$y$ plane, and the gear’s initial orientation is randomized within $45^{\circ}$ of its nominal pose. In this task, the gripper remains closed to maintain a stable grasp on the gear during precise alignment and placement.
Appendix C Dataset Details
@@
We collect expert demonstrations using a 3Dconnexion SpaceMouse for all tasks in the ManiFeel benchmark, including six simulated tasks and one real-world task under two lighting conditions. Table 5 summarizes the amount of demonstrations, camera views, tactile sensors, and action dimensions (ActD) recorded per task. Tasks with an action dimension (ActD) of 7 include gripper open/close actions, while those with an ActD of 6 use a fixed gripper state, which is excluded from the action space.
Each simulation task is recorded with three RGB camera views (front, side, wrist) and two tactile sensors (left and right fingers), whereas the real-world setup includes a front-facing Intel RealSense camera and a single GelSight tactile sensor on the left finger. While not all camera views and tactile sensors are required as inputs during policy training, our dataset captures diverse sensor observations to support flexible usage in future research. The specific observation setup used for ManiFeel policy training is described in Appendix F.
The demonstrations are recorded at 101010\,10Hz and include synchronized visual, tactile, and proprioceptive observations. The proprioceptive feedback includes the end-effector’s position and orientation, resulting in a 7-dimensional state vector, where orientation is represented using a quaternion.
@@
Table 5: Dataset Summary. Cam: number of camera viewpoints recorded in the dataset; Tac: number of tactile sensors recorded; PropD: proprioceptive feedback dimension; ActD: action dimension; Demo: number of human demonstrations collected.
@@
Tactile Sorting 3 2 7 7 51
Tactile Exploration 3 2 7 7 41
Blind Insertion 3 2 7 6 51
USB Insertion 3 2 7 6 61
Power Plug Insertion 3 2 7 6 101
Gear Assembly 3 2 7 6 21
@@
Tactile Sorting (Normal) 1 1 7 5 103
Tactile Sorting (Dim) 1 1 7 5 80
Representative examples of the demonstration trajectories are visualized in Fig. 3. In these scenarios, where visual input alone is often insufficient due to occlusions or degraded lighting, operators rely on real-time tactile images displayed on a monitor to guide precise behaviors, such as identifying surface properties for object classification or estimating in-hand pose for alignment. For instance, in the tactile sorting task under dim lighting, human operators struggle to distinguish between visually similar objects like a table tennis ball and a golf ball. Tactile feedback enables operators to detect surface texture differences and correct initial grasping errors. These observations highlight the importance of tactile input not only for effective teleoperation but also for collecting high-quality data to support supervised visuotactile policy learning. All of the collected datasets will be open-sourced to support future research in supervised visuotactile policy learning.
(a) Tactile sorting
(b) Tactile exploration
Figure 3: Samples of teleoperated demonstration data in simulation for two representative tasks: (a) tactile sorting and (b) tactile exploration, highlighting scenarios with severely degraded or occluded visual input.
Appendix D Pipeline Details
D.1 Visual Encoder
We adopt a ResNet18 CNN encoder (without pretraining) with two key modifications as described in [1]: (1) the global average pooling layer is replaced with a spatial softmax pooling layer to retain spatial information; and (2) Batch Normalization is replaced with Group Normalization to ensure stable training. Each camera view is processed using a separate encoder. For visuotactile policies that do not rely on pretrained tactile representations, the same ResNet18 encoder is also used to extract features from tactile images.
D.2 Tactile Representations
@@
AnyTouch [23] is also a tactile representation built upon vision transformer backbone. It is pre-trained on dataset that includes aligned tactile images and videos collected from four different visuo-tactile sensors. AnyTouch jointly leverages pixel-level reconstruction and multi-sensor alignment objectives to learn semantic features. It enables robust perception and effective cross-sensor transfer, demonstrating strong generalization across both static and dynamic tasks. We use the checkpoint released by AnyTouch [23] for both simulation and real-world experiments. The model’s encoder is based on a ViT-Large backbone, with a total of 305 M parameters. As shown in Table 8, the large size of AnyTouch significantly slows down training for policies, reducing its flexibility.
D.3 Policies
@@
Appendix E Training Details
Table 6: Hyperparameters for training and inference. To: observation horizon; Ta: action horizon; Tp: action prediction horizon; ImgRes: environment RGB image resolution; TacImgRes: tactile image resolution; Lr: learning rate; DiffT.: number of training diffusion iterations; DiffE.: number of inference diffusion iterations.
H-Param To Ta Tp ImgRes TacImgRes Lr DiffT DiffE
Tactile Sorting 2 8 16 256x256 256x256 1e-4 100 100
Tactile Exploration 2 8 16 256x256 256x256 1e-4 100 100
Blind Insertion 2 8 16 256x256 256x256 1e-4 100 100
USB Insertion 2 8 16 256x256 256x256 1e-4 100 100
Power Plug Insertion 2 8 16 256x256 256x256 1e-4 100 100
Gear Assembly 2 8 16 256x256 256x256 1e-4 100 100
Real Sorting (Normal) 2 6 16 320x240 320x240 1e-4 75 16
Real Sorting (Dim) 2 8 16 320x240 320x240 1e-4 75 16
We summarize the training and inference hyperparameters in Table 6, which are consistent across all policies. The original visual image resolution is $256{\times}256$, while the tactile images are captured at $240{\times}320$. For training, we consistently resize tactile images to $256{\times}256$ when using the ResNet18 tactile encoder and UniT tactile representation, while visual inputs are maintained at $256{\times}256$ for the visual encoder. For T3 and AnyTouch, we use $224{\times}224$ tactile image inputs to match the ViT backbone architecture. For Diffusion Policy, we adopt DDPM for simulation experiments and switch to DDIM during real-world experiments to enable faster inference. For Equivariant Diffusion Policy, we use DDIM.
Appendix F Additional Details on Experiments
F.1 Task Setup
Table 7 summarizes the task configurations used for policy training and evaluation in both simulation and real-world experiments, including the number of cameras, tactile sensors, and training datasets per task. In all experiments, only tactile images from the left finger sensor are used. Tasks with an action dimension (ActD) of 7 include gripper open/close actions, while those with an ActD of 6 use a fixed gripper state, which is excluded from the action space. For real-world experiments, we use an OSC-yaw controller that controls end-effector position and yaw rotation, while roll and pitch are kept fixed throughout the execution. All code and configuration files for training and rollout will be open-sourced to facilitate reproducibility of the results.
Table 7: Tasks Setup in the Experiments. Ctrl: control type; Cam: number of cameras; Tac: number of tactile sensors; ActD: action dimension; Demo: number of demonstrations used for training; Lighting: lighting condition; Occlusion: whether the camera view is occluded.
@@ F.1 Task Setup
Simulation experiment
Tactile Sorting Relative Pose 1 (Front) 1 7 50 Dim No
Tactile Exploration Relative Pose 1 (Front) 1 7 40 Normal Yes
Blind Insertion Relative Pose 1 (Front) 1 6 30 Normal Yes
USB Insertion Relative Pose 1 (Wrist) 1 6 60 Normal No
Power Plug Insertion Relative Pose 1 (Wrist) 1 6 90 Normal No
Gear Assembly Relative Pose 1 (Wrist) 1 6 20 Normal No
Real-world experiment
Tactile Sorting (normal) Relative Pose 1 1 5 103 Normal No
Tactile Sorting (dim) Relative Pose 1 1 5 80 Dim No
F.2 Learning Curves
In this section, we present learning curves that show success rates over training epochs for all experiments reported in Section 3. Each curve represents the mean success rate across three random seeds (solid line), and the shaded region denotes the range between the maximum and minimum values at each epoch. Policies are evaluated every 10 epochs using 50 different environment initializations to ensure statistical robustness. Figure 4 presents the learning curves from the modality benchmarking experiment (Section 3.1), comparing vision-only, tactile-only, and visuotactile inputs across all six ManiFeel tasks. For tactile sorting, tactile exploration, and blind insertion, we additionally evaluate a state-only policy to analyze the contribution of proprioceptive feedback. The state-only input refers to the 7-dimensional proprioceptive feedback of the end-effector pose (position and orientation), which is also included in all other sensing modality configurations. Notably, for tactile exploration and blind insertion, where visual input is severely occluded, the vision-only and state-only policies perform similarly, suggesting that the performance in vision-only settings is largely driven by proprioceptive input rather than visual cues. Figs. 5 and 6 present learning curves for the tactile representation (Section 3.2) and policy benchmarking (Section 3.3) experiments, respectively, across three representative tasks: tactile sorting, tactile exploration, and gear assembly.
Tactile Sorting
Tactile Exploration
Blind Insertion
USB Insertion
Power Plug Insertion
Gear Assembly
Figure 4: Leanring curves comparing sensing modalities across six manipulation tasks. Each plot shows success rates over training epochs. The first row includes tactile sorting, tactile exploration, and blind insertion, where all four sensory configurations are evaluated: visuotactile, tactile-only, vision-only, and state-only. The second row includes USB insertion, power plug insertion, and gear assembly, where only visuotactile, tactile-only, and vision-only configurations are evaluated. Note that all visuotactile, tactile-only, and vision-only policies also include proprioceptive state inputs.
Tactile Sorting
Tactile Exploration
Gear Assembly
Figure 5: Learning curves across tactile-augmented policies using different tactile representations, for three representative tasks, including tactile sorting, tactile exploration, and gear assembly. For the tactile exploration task, the policy uses only tactile observations without visual input.
Tactile Sorting (Vision-only)
Tactile Exploration (Vision-only)
Gear assembly (Vision-only)
Tactile Sorting (Visuotactile)
Tactile Exploration (Tactile-only)
Gear assembly (Visuotactile)
Figure 6: Learning curves for three representative tasks, including tactile sorting, tactile exploration, and gear assembly, evaluated across different policy architectures. Top: vision-only policy. Bottom: tactile-augmented policy, where visual input is included for tactile sorting and gear assembly, but excluded for tactile exploration.
F.3 Runtime Analysis of Tactile Representations

### v2

@@ Appendix A Additional Details on Task Environments
Insertion Tasks (Sim). For the peg, USB, power plug, and gear assembly tasks, both the socket location and the in-hand orientation of the object (peg, USB, plug, or insertion gear) are randomized at the start of each rollout. The socket position or gear base is sampled within a (5-5 cm, +5+5 cm) range along the $x$ and $y$ axes, and the object’s in-hand orientation is randomized within $10^{\circ}$ of its nominal pose. The gripper remains closed throughout to maintain a stable grasp during precise alignment and insertion (see Fig. 3).
Screwing Tasks (Sim). This category includes nut–bolt assembly and bulb installation. For both tasks, the bolt or socket position is randomized within (5-5 cm, +5+5 cm) along the $x$ and $y$ axes from its nominal location at the start of each rollout, while the in-hand rotation of the object (nut or bulb) is fixed initially. The robot end-effector’s roll and pitch angles are also fixed during execution. In this setup, the robot hand first approaches the bolt or socket and then performs the screwing motion. At this stage, the gripper’s open–close action and applied grasping force are regulated to maintain stable contact and ensure reliable threading.
Peg Reorientation (Sim). The robot must perform self-alignment by leveraging external contact with the socket rim to rotate the peg in hand before insertion into a socket placed inside an enclosed container, emulating how humans manipulate or orient objects in confined spaces without visual access. The socket position remains fixed within a box, while the peg’s initial in-hand orientation is randomized up to $30^{\circ}$. The gripper remains closed throughout to ensure a stable grasp during in-hand reorientation and insertion (see Fig. 4a).
Object Search (Sim). This task requires the robot to pick up a cube-shaped battery from inside a closed container, emulating how humans retrieve objects from a bag without visual access. At the beginning of each demonstration or evaluation, the object’s position is randomly selected between two predefined locations along the $x$-axis within the container. Because the interior is fully occluded, the robot relies solely on tactile feedback to explore, localize, and grasp the object (see Fig. 4b).
Ball Sorting (Sim). This task requires the robot to pick up the golf ball from a pair of golf and table-tennis balls placed under dim lighting conditions to simulate degraded visual perception. A golf ball and a table tennis ball are placed on a ball stand whose position is randomized within (5-5 cm, +5+5 cm) along the $x$ and $y$ axes from its nominal position. The ball positions (golf vs. table tennis) are swapped across episodes to increase task diversity (see Fig. 4c).
Gear Assembly (Real). In the real-world setting, the robot first grasps the insertion gear and then aligns and mounts it onto the central shaft of a base featuring two side gears. During data collection and evaluation rollouts, the gear base is slightly rotated in-plane from its nominal orientation to introduce variability (see Fig. 10a).
Bulb Installation (Real). In this setup, the bulb is initially placed in its socket, and the robot moves to grasp and screw it into place. The initial screw depth of the bulb within the socket is randomized across rollouts to introduce variability, resulting in uncertainty in the number of turns required for proper tightening (see Fig. 10b).
Ball Sorting (Real). The robot must classify and sort two visually similar spherical objects, a golf ball and a table tennis ball, into orange and blue bowls, respectively. The task is conducted under normal and dim lighting conditions. The positions of the two balls are swapped between rollouts to introduce variation (see Figs. 10d10c).
Appendix B Dataset Details
@@
We collect expert demonstrations using a 3Dconnexion SpaceMouse for all tasks in the ManiFeel benchmark, including $9$ simulation and $4$ real-world tasks. Table V summarizes the amount of demonstrations, camera views, tactile sensors, and action dimensions (ActD) recorded per task. Tasks with an action dimension of $7$ include gripper open/close actions, while those with an ActD of $6$ use a fixed gripper state, which is excluded from the action space. In the nut–bolt assembly, bulb installation, and object search tasks, although our dataset records ActD of $7$ for pipeline consistency, the end-effector rotations are constrained during execution, where roll–pitch angles are fixed for the nut–bolt assembly and bulb installation, and roll–pitch–yaw angles are fixed for the object-search task. For real-world experiments, the roll and pitch angles are fixed throughout execution, resulting in an ActD of $5$.
@@
TABLE V: Dataset Summary. Cam: number of camera viewpoints recorded in the dataset; Tac: number of tactile sensors recorded; PropD: proprioceptive feedback dimension; ActD: action dimension; Demo: number of human demonstrations collected.
@@
Peg Insertion 3 2 7 6 50
USB Insertion 3 2 7 6 50
Power Plug Insertion 3 2 7 6 50
Gear Assembly 3 2 7 6 100
Nut–bolt Assembly 3 2 7 7 30
Bulb Installation 3 2 7 7 20
Peg Reorientation 3 2 7 6 50
Object Search 3 2 7 7 50
Ball Sorting (Normal) 3 2 7 7 80
Ball Sorting (Dim) 3 2 7 7 80
@@
Gear Assembly 1 1 7 5 80
Bulb Installation 1 1 7 5 50
Ball Sorting (Normal) 1 1 7 5 80
Ball Sorting (Dim) 1 1 7 5 80
Each simulation task is recorded with three RGB camera views (front, side, wrist) and two tactile sensors (left and right fingers), whereas the real-world setup includes a front-facing Intel RealSense camera and a single GelSight R1.5 tactile sensor on the right finger. The demonstrations are recorded at 1010\,Hz and include synchronized visual, tactile, and proprioceptive observations. The proprioceptive feedback includes the end-effector’s position and orientation, resulting in a 7-dimensional state vector, where orientation is represented using a quaternion. While not all camera views and tactile sensors are required as inputs during policy training, our dataset captures diverse sensor observations to support flexible use in future research. The specific observation setup used for our experiments is described in Appendix D.
Appendix C Pipeline Details
C-A Visual and Tactile Encoders
We adopt a ResNet18 CNN encoder (without pretraining) with two key modifications as described in [10]: (1) the global average pooling layer is replaced with a spatial softmax pooling layer to retain spatial information; and (2) Batch Normalization is replaced with Group Normalization to ensure stable training. Each camera view is processed using a separate encoder. For visuotactile policies without pretrained tactile representations, we use the same ResNet18 encoder to extract latent embeddings from both tactile images and tactile force fields.
C-B Tactile Representations
@@
AnyTouch [14] is also a tactile representation built upon vision transformer backbone. It is pre-trained on dataset that includes aligned tactile images and videos collected from four different visuo-tactile sensors. AnyTouch jointly leverages pixel-level reconstruction and multi-sensor alignment objectives to learn semantic features. It enables robust perception and effective cross-sensor transfer, demonstrating strong generalization across both static and dynamic tasks. We use the checkpoint released by AnyTouch [14] for both simulation and real-world experiments. The model’s encoder is based on a ViT-Large backbone, with a total of 305 M parameters. As shown in Table VIII, the large size of AnyTouch significantly slows down training for policies, reducing its flexibility.
C-C Policies
@@
Appendix D Additional Details on Experiments
D-A Experimental Setup
Table VI summarizes the task configurations used for policy training and evaluation in both simulation and real-world experiments, including the number of cameras, tactile sensors, and training demonstrations per task. Two simulation setups are used in this work. The first setup, corresponding to the results reported in Sections IV-A – IV-C, is used to benchmark the proposed ManiFeel task suite and analyze the contribution of different sensing modalities, tactile representations, and policy architectures. The second simulation setup mirrors the number of demonstrations and conditions of the real-world experiments (Section IV-D) to enable a direct comparison between simulated and physical environments. In all experiments, only tactile signals from the right finger sensor are used. For real-world experiments, we employ an OSC-yaw controller that regulates the end-effector position and yaw rotation while keeping roll and pitch fixed throughout execution.
TABLE VI: Tasks setup in the experiments. Ctrl: control type; Cam: number of cameras; Tac: number of tactile sensors; ActD: action dimension; Demo: number of demonstrations used for training; Lighting: lighting condition; Occlusion: whether the camera view is occluded.
@@ D-A Experimental Setup
Simulation Experiment (Sections IV-A – IV-C)
Peg Insertion Relative Pose 1 (Wrist) 1 6 50 Normal No
USB Insertion Relative Pose 1 (Wrist) 1 6 50 Normal No
Power Plug Insertion Relative Pose 1 (Wrist) 1 6 50 Normal No
Gear Assembly Relative Pose 1 (Wrist) 1 6 50 Normal No
Nut–bolt Assembly Relative Pose 1 (Wrist) 1 6 30 Normal No
Bulb Installation Relative Pose 1 (Wrist) 1 6 20 Normal No
Peg Reorientation Relative Pose 1 (Front) 1 6 50 Normal Yes
Object Search Relative Pose 1 (Front) 1 7 50 Normal Yes
Ball Sorting (Dim) Relative Pose 1 (Front) 1 7 50 Dim No
Simulation Experiment (Section IV-D)
Gear Assembly Relative Pose 1 (Wrist) 1 6 80 Normal No
Bulb Installation Relative Pose 1 (Wrist) 1 5 20 Normal No
Ball Sorting (Normal) Relative Pose 1 (Front) 1 7 80 Normal No
Ball Sorting (Dim) Relative Pose 1 (Front) 1 7 80 Dim No
Real-world Experiment (Section IV-D)
Gear Assembly Relative Pose 1 (Wrist) 1 5 80 Normal No
Bulb Installation Relative Pose 1 (Wrist) 1 5 20 Normal No
Ball Sorting (Normal) Relative Pose 1 (Front) 1 5 80 Normal No
Ball Sorting (Dim) Relative Pose 1 (Front) 1 5 80 Dim No
Figure 12: Success rate of gear assembly with increasing demonstrations across sensing configurations. Performance saturates around $80$ demos, indicating a balance point between data efficiency and policy performance.
To determine an appropriate number of demonstrations for training, we conducted a scaling study on the gear assembly task, as illustrated in Fig. 12. The results show that success rates increase steadily with the number of demonstrations but begin to saturate at around $80$ demonstrations across all sensing configurations. Therefore, we adopt $80$ demonstrations as the shared training set size for both simulation and real-world comparative experiments on gear assembly and ball sorting tasks, balancing data efficiency and policy stability. For the bulb installation task, we use $20$ demonstrations due to the inherently longer teleoperation trajectories, which result in substantially heavier datasets and higher training costs. Despite the smaller dataset, the policy achieves a high success rate (over $70\%$) in simulation, indicating that performance has already converged within this data regime. This choice thus represents a practical trade-off between data collection efficiency and policy performance.
All code, datasets, and configuration files used for training and rollout will be released publicly to facilitate full reproducibility and extension of the ManiFeel benchmark.
D-B Training Details
We summarize the training and inference hyperparameters in Table VII, which are consistent across all policies. The original visual image resolution is $256{\times}256$, while the tactile images and tactile force fields are recorded at $240{\times}320$ and $10{\times}14$, respectively. For training, we consistently resize tactile images to $256{\times}256$ when using the ResNet18 tactile encoder and UniT tactile representation, while visual inputs are maintained at $256{\times}256$ for the visual encoder. For T3 and AnyTouch, we use $224{\times}224$ tactile image inputs to match the ViT backbone architecture. For Diffusion Policy, we adopt DDPM for simulation experiments and switch to DDIM during real-world experiments to enable faster inference. For Equivariant Diffusion Policy (EquiDP), we use DDIM.
TABLE VII: Hyperparameters for training and inference. To: observation horizon; Ta: action horizon; Tp: action prediction horizon; ImgRes: environment RGB image resolution; TacImgRes: tactile image resolution; TacFFRes: tactile force-field resolution; Lr: learning rate; DiffT.: number of training diffusion iterations; DiffE.: number of inference diffusion iterations.
H-Param To Ta Tp ImgRes TacImgRes TacFFRes Lr DiffT DiffE
Simulation
Peg Insertion 2 8 16 256×256 240×320 10×14 8e-5 100 100
USB Insertion 2 8 16 256×256 240×320 10×14 8e-5 100 100
Power Plug Insertion 2 8 16 256×256 240×320 10×14 8e-5 100 100
Gear Assembly 2 8 16 256×256 240×320 10×14 8e-5 100 100
Nut–bolt Assembly 2 8 16 256×256 240×320 10×14 8e-5 100 100
Bulb Installation 2 8 16 256×256 240×320 10×14 8e-5 100 100
Peg Reorientation 2 8 16 256×256 240×320 10×14 8e-5 100 100
Object Search 2 8 16 256×256 240×320 10×14 8e-5 100 100
Ball Sorting 2 8 16 256×256 240×320 10×14 8e-5 100 100
Real-World
Gear Assembly 2 6 16 320×240 160×240 10×14 8e-5 75 16
Bulb Installation 2 8 16 320×240 160×240 10×14 8e-5 75 16
Ball Sorting (Normal) 2 6 16 320×240 160×240 10×14 8e-5 75 16
Ball Sorting (Dim) 2 8 16 320×240 160×240 10×14 8e-5 75 16
D-C Runtime Analysis of Tactile Representations